Dans dad is 28 years older than Dan. The sum of their ages is 64. How old is Dan?
Answer by benazir.sj@gmail.com(109)

Let dans age =x
dans dads age =x+28
sum of their ages =64
x+x+28 =64
2x=64-28
2x=36
x=18
Dan is 18 years and
his dad is 46 years old